hi. im having a really hard time solving a logarithmic equation. heres the first problem log4 x=3/2 can you explain it to me?
Answer by nerdybill(5404)

Apply "log rules"
You can review them at:
http://www.purplemath.com/modules/logrules.htm
.
log4 x=3/2
x = 4^(3/2)
x = sqrt(4^3)
x = sqrt(64)
x = +- 8
.
We can toss out the negative solution because you can't take a log of a negative number.
Therefore,
x = 8